### v1

Abstract: The geometric classifications of complex $4$-dimensional nilpotent Lie-Yamaguti algebras, $4$-dimensional nilpotent Bol algebras, and $4$-dimensional nilpotent compatible Lie algebras are given.
Keywords: Lie-Yamaguti algebra, Bol algebra, compatible Lie algebra, nilpotent algebra, rigid algebra, irreducible components, geometric classification.
MSC2020: 17A30 (primary); 17B63, 14L30 (secondary).
Introduction
The study of algebras endowed with two or more multiplication operations is not a recent development; nevertheless, it continues to attract considerable attention within the mathematical community due to its theoretical richness and potential applications. [27, 6, 7, 10, 40]. The class of Poisson algebras is the most well-known example of algebras featuring two multiplication operations.
The notion of the Lie triple system was formally introduced as an algebraic object by Jacobson in connection with problems arising from quantum mechanics. The example of a Lie triple system that arose from quantum mechanics is known as the Meson field, as described by Jacobson [21]. The notion of Lie-Yamaguti algebras is a generalization of Lie triple systems and Lie algebras, derived from Nomizu’s work on the invariant affine connections on homogeneous spaces in the 1950s [34]. The notion of a Lie-Yamaguti algebra is a natural abstraction made by Yamaguti of Nomizu’s considerations. Yamaguti called these systems general Lie triple systems [41]. Kikkawa renamed the notion of a general Lie triple system as Lie triple algebra and showed how to “integrate” these algebras to nonassociative multiplications on reductive homogeneous spaces [24]. Kinyon and Weinstein observed that Lie triple algebras, which they called “Lie-Yamaguti algebras” in their paper, can be constructed from Leibniz algebras [25]. The interplay between their binary and ternary operations reflects the geometric and algebraic intricacies of the spaces from which they originate, making Lie-Yamaguti algebras a significant object of study in both mathematics and theoretical physics. Lie-Yamaguti algebras have been studied by several authors in [38, 39, 18, 17, 42, 41, 1]. In particular, Sagle constructed some remarkable examples of Lie-Yamaguti algebras arising from reductive homogeneous spaces in differential geometry. In [42], the deformation and extension theory of Lie-Yamaguti algebras is studied, and the notion of Nijenhuis operators for Lie-Yamaguti algebras is introduced to describe trivial deformations. The algebraic classification of the nilpotent Lie-Yamaguti algebras up to dimension four was given in [1].
[…]
A compatible Lie algebra is a pair of Lie algebras such that any linear combination of these two Lie brackets is still a Lie bracket. Such structures are closely related to the (infinitesimal) deformations of Lie algebras. Compatible Lie algebras are considered in many fields in mathematics and mathematical physics, such as the study of the classical Yang-Baxter equation [14], integrable equations of the principal chiral model type [15], elliptic theta functions [35], and loop algebras over Lie algebras [16]. Compatible Lie algebras have a natural one-to-one correspondence to compatible linear Poisson structures, whereas the latter are important Poisson structures related to bi-Hamiltonian structures. Recall that a bi-Hamiltonian structure is a pair of Poisson structures on a manifold which are compatible, i.e., their sum is again a Poisson structure [26].
Degenerations and geometric properties of a variety of algebras have been an object of study since 1970’s. Gabriel [13] described the irreducible components of the variety of $4$-dimensional unital associative algebras. Mazzola [32] classified algebraically and geometrically the variety of unital associative algebras of dimension $5$. Burde and Steinhoff [9] constructed the graphs of degenerations for the varieties of $3$-dimensional and $4$-dimensional Lie algebras over ${\mathbb{C}}$. Grunewald and O’Halloran [19] calculated the degenerations for the nilpotent Lie algebras of dimension up to $5$.
The study of degenerations of algebras is very rich and closely related to deformation theory. It offers an insightful geometric perspective on the subject and has been the object of a lot of research. One of the main problems of the geometric classification of a variety of algebras is the description of its irreducible components. In the case of finitely many orbits (i.e., isomorphism classes), the irreducible components are determined by the rigid algebras – algebras whose orbit closure is an irreducible component of the variety under consideration. There are many results related to the deformation and geometric classification of algebras in different varieties of associative and non-associative algebras, superalgebras and Poisson algebras see [3, 4, 9, 7, 30, 11, 22, 31, 23] and references in [22, 23].
In the present paper, we continue the study of the geometric classification of algebras and describe all irreducible components of the variety of four-dimensional nilpotent Lie-Yamaguti algebras, four-dimensional nilpotent Bol algebras, three and four-dimensional nilpotent compatible Lie algebras. Our main results are summarized below.
Theorem A.
The variety of complex four-dimensional nilpotent Lie-Yamaguti algebras has dimension $14$. It is defined by one rigid algebra and two $1$-parametric families of algebras and can be described as the closure of the union of $\mathrm{GL}_{4}(\mathbb{C})$-orbits of the algebras given in Theorem 18.
Theorem B.
The variety of complex four-dimensional nilpotent Bol algebras has dimension $15$. It is defined by one rigid algebra and one $1$-parametric family of algebras and can be described as the closure of the union of $\mathrm{GL}_{4}(\mathbb{C})$-orbits of the algebras given in Theorem 19.
Theorem C.
The variety of complex four-dimensional nilpotent compatible Lie algebras has dimension $13$. It is defined by one $1$-parametric family of algebras and can be described as the closure of the union of $\mathrm{GL}_{4}(\mathbb{C})$-orbits of the algebras given in Theorem 21.
1 Preliminaries: the algebraic classification
All the algebras below will be over $\mathbb{C}$ and all the linear maps will be $\mathbb{C}$-linear. For simplicity, every time we write the multiplication table of an algebra the products of basic elements whose values are zero or can be recovered from the commutativity or from the anticommutativity are omitted. The notion of a nontrivial algebra means that the multiplication is nonzero. In this section, we present the necessary concepts related to Lie-Yamaguti algebras, Bol algebras, and compatible Lie algebras. We present the algebraic classification of four-dimensional nilpotent Lie-Yamaguti algebras, four-dimensional nilpotent Bol algebras.
Definition 1.
[…]
${\mathfrak{D}}[x,y]=[{\mathfrak{D}}(x),y]+[x,{\mathfrak{D}}(y)],\ {\mathfrak{D}}[x,y,z]=[{\mathfrak{D}}(x),y,z]+[x,{\mathfrak{D}}(y),z]+[x,y,{\mathfrak{D}}(z)].$
[…]
We denote the set of all derivations of the algebra $\mathcal{L}$ by $\mathfrak{Der}(\mathcal{L})$.
Definition 2.
[…]
$\mathcal{L}^{(1)}:=\mathcal{L},\quad\mathcal{L}^{(n)}:=\sum\limits_{i+j=n}[\mathcal{L}^{(i)},\mathcal{L}^{(j)}]+\sum\limits_{i+j+k=n+1}[\mathcal{L}^{(i)},\mathcal{L}^{(j)},\mathcal{L}^{(k)}],\ i,j,k\geq 1,n\geq 2.$
[…]
Definition 3 (see [25]).
A Lie-Yamaguti algebra is a vector space $\mathcal{L}$ with a bilinear multiplication [,][-,-][ - , - ] and a trilinear multiplication [,,][-,-,-][ - , - , - ] satisfying:
$\begin{array}[]{ll}({\rm LY1})&[x,y]+[y,x]=0,\\ ({\rm LY2})&[x,y,z]+[y,x,z]=0,\\ ({\rm LY3})&[x,y,z]+[y,z,x]+[z,x,y]+[[x,y],z]+[[y,z],x]+[[z,x],y]=0,\\ ({\rm LY4})&[[x,y],z,u]+[[y,z],x,u]+[[z,x],y,u]=0,\\ ({\rm LY5})&[x,y,[u,v]]-[[x,y,u],v]-[u,[x,y,v]]=0,\\ ({\rm LY6})&[u,v,[x,y,z]]-[[u,v,x],y,z]-[x,[u,v,y],z]-[x,y,[u,v,z]]=0.\\ \end{array}$
Example 4.
The Lie-Yamaguti algebra with $[x,y]=0$ for any $x,y\in\mathcal{L}$ is exactly a Lie triple system, closely related to symmetric spaces, while the Lie-Yamaguti algebra with $[x,y,z]=0$ for any $x,y,z\in\mathcal{L}$ is the Lie algebra.
Theorem 5 (see [1]).
Let $\mathcal{L}$ be a $4$-dimensional complex nilpotent Lie-Yamaguti algebra. Then $\mathcal{L}$ is isomorphic to one of the following pairwise non-isomorphic algebras:
[…]
$\mathcal{L}_{02}$ : $[e_{1},e_{2},e_{1}]=e_{3}.$
$\mathcal{L}_{03}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{2},e_{1}]=e_{3}.$
$\mathcal{L}_{04}$ :::: $[e_{1},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{05}$ :::: $[e_{1},e_{2}]=e_{4},$ $[e_{2},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{06}$ :::: $[e_{2},e_{3},e_{2}]=e_{4},$ $[e_{3},e_{1},e_{3}]=e_{4}$.
$\mathcal{L}_{07}$ :::: $[e_{1},e_{2}]=e_{4},$ $[e_{2},e_{3},e_{2}]=e_{4},$ $[e_{3},e_{1},e_{3}]=e_{4}.$
$\mathcal{L}_{08}$ :::: $[e_{1},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{2}]=e_{4},$ $[e_{3},e_{1},e_{3}]=e_{4}.$
$\mathcal{L}_{09}$ :::: $[e_{2},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{2}]=e_{4},$ $[e_{3},e_{1},e_{3}]=e_{4}.$
$\mathcal{L}_{10}$ :::: $[e_{2},e_{3},e_{1}]=e_{4},$ $[e_{3},e_{1},e_{2}]=e_{4},$ $[e_{2},e_{3},e_{2}]=e_{4}.$ $[e_{2},e_{1},e_{3}]=2e_{4}.$
$\mathcal{L}_{11}$ :::: $[e_{1},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{1}]=e_{4},$ $[e_{3},e_{1},e_{2}]=e_{4},$
$[e_{2},e_{3},e_{2}]=e_{4}.$ $[e_{2},e_{1},e_{3}]=2e_{4}.$
$\mathcal{L}_{12}$ :::: $[e_{2},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{1}]=e_{4},$ $[e_{3},e_{1},e_{2}]=e_{4},$
$[e_{2},e_{3},e_{2}]=e_{4}.$ $[e_{2},e_{1},e_{3}]=2e_{4}.$
$\mathcal{L}_{13}$ :::: $[e_{1},e_{2}]=e_{4},$ $[e_{2},e_{3},e_{1}]=e_{4},$ $[e_{3},e_{1},e_{2}]=e_{4},$
$[e_{2},e_{3},e_{2}]=e_{4}.$ $[e_{2},e_{1},e_{3}]=2e_{4}.$
$\mathcal{L}_{14}$ :::: $[e_{1},e_{2}]=e_{4},$ $[e_{2},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{1}]=e_{4},$ $[e_{3},e_{1},e_{2}]=e_{4},$
$[e_{2},e_{3},e_{2}]=e_{4}.$ $[e_{2},e_{1},e_{3}]=2e_{4}.$
$\mathcal{L}_{15}$ :::: $[e_{1},e_{2}]=e_{4},$ $[e_{1},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{1}]=e_{4},$ $[e_{3},e_{1},e_{2}]=e_{4},$
$[e_{2},e_{3},e_{2}]=e_{4},$ $[e_{2},e_{1},e_{3}]=2e_{4}.$
$\mathcal{L}_{16}^{\alpha}$ :::: $[e_{2},e_{3},e_{1}]=\alpha e_{4},$ $[e_{1},e_{2},e_{3}]=-(1+\alpha)e_{4},$ $[e_{3},e_{1},e_{2}]=e_{4}.$
$\mathcal{L}_{17}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{4},$ $[e_{1},e_{2},e_{3}]=-(1+\alpha)e_{4},$ $[e_{2},e_{3},e_{1}]=\alpha e_{4},$ $[e_{3},e_{1},e_{2}]=e_{4}.$
$\mathcal{L}_{18}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{4},$ $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{3}]=-(1+\alpha)e_{4},$
$[e_{2},e_{3},e_{1}]=\alpha e_{4},$ $[e_{3},e_{1},e_{2}]=e_{4}.$
$\mathcal{L}_{19}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{4},$ $[e_{1},e_{3}]=e_{4},$ $[e_{2},e_{3}]=e_{4},$ $[e_{3},e_{1},e_{2}]=e_{4}.$
$[e_{2},e_{3},e_{1}]=\alpha e_{4},$ $[e_{1},e_{2},e_{3}]=-(1+\alpha)e_{4},$
$\mathcal{L}_{20}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=\alpha e_{4}.$
$\mathcal{L}_{21}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{2},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{4}.$
$\mathcal{L}_{22}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{2},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{23}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=\alpha e_{4},$ $[e_{1},e_{2},e_{1}]=e_{4},$ $[e_{2},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{24}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{2},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{4},$ $[e_{2},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{25}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{26}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{2},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{27}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3},e_{2}]=e_{4},$ $[e_{2},e_{3},e_{1}]=e_{4}.$
$\mathcal{L}_{28}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{3},e_{2}]=e_{4},$ $[e_{2},e_{3},e_{1}]=e_{4}.$
$\mathcal{L}_{29}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=e_{4},$ $[e_{2},e_{3}]=e_{4},$ $[e_{1},e_{3},e_{2}]=e_{4},$
$[e_{2},e_{3},e_{1}]=e_{4}.$
$\mathcal{L}_{30}$ :::: $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3}.$
$\mathcal{L}_{31}$ :::: $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{2},e_{3}]=e_{4},$ $[e_{1},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{32}$ :::: $[e_{1},e_{2}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{2},e_{3}]=e_{4},$ $[e_{1},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{33}$ :::: $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{2},e_{3}]=e_{4},$ $[e_{1},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{34}$ :::: $[e_{1},e_{2}]=e_{4},$ $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{2},e_{3}]=e_{4},$
$[e_{1},e_{3},e_{2}]=e_{4}.$
$\mathcal{L}_{35}$ :::: $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{3},e_{1}]=e_{4},$ $[e_{1},e_{2},e_{2}]=e_{4}.$
$\mathcal{L}_{36}$ :::: $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{3},e_{1}]=e_{4},$ $[e_{1},e_{2},e_{2}]=e_{4}.$
$\mathcal{L}_{37}$ :::: $[e_{1},e_{2}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{3},e_{1}]=e_{4},$ $[e_{1},e_{2},e_{2}]=e_{4}.$
$\mathcal{L}_{38}$ :::: $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{3},e_{1}]=e_{4}.$
$\mathcal{L}_{39}$ :::: $[e_{1},e_{2}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{3},e_{1}]=e_{4}.$
$\mathcal{L}_{40}$ :::: $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{3},e_{1}]=e_{4}.$
$\mathcal{L}_{41}$ :::: $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{2},e_{2}]=e_{4}.$
$\mathcal{L}_{42}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{2},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{2},e_{3}]=-e_{4},$
$[e_{1},e_{3},e_{2}]=-e_{4}.$
$\mathcal{L}_{43}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=e_{4},$ $[e_{2},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$
$[e_{1},e_{2},e_{3}]=-e_{4},$ $[e_{1},e_{3},e_{2}]=-e_{4}.$
$\mathcal{L}_{44}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{3}+e_{4},$ $[e_{1},e_{3}]=\alpha e_{4},$ $[e_{2},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$
$[e_{1},e_{2},e_{3}]=-e_{4},$ $[e_{1},e_{3},e_{2}]=-e_{4}.$
$\mathcal{L}_{45}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{3},e_{1}]=\alpha e_{4}.$
$\mathcal{L}_{46}^{\alpha\neq 0}$ :::: $[e_{1},e_{2}]=e_{3}+\alpha e_{4},$ $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{3},e_{1}]=e_{4}.$
$\mathcal{L}_{47}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{2},e_{2}]=e_{4},$
$[e_{1},e_{3},e_{1}]=\alpha e_{4}.$
$\mathcal{L}_{48}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{3},e_{1}]=e_{4}.$
$\mathcal{L}_{49}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{2},e_{2}]=e_{4},$ $[e_{1},e_{3},e_{1}]=e_{4}.$
$\mathcal{L}_{50}$ :::: $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{2},e_{2}]=e_{4}.$
$\mathcal{L}_{51}$ :::: $[e_{1},e_{2}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3}.$
$\mathcal{L}_{52}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{2},e_{2}]=e_{4}.$
Where $\mathcal{L}_{46}^{0}\cong\mathcal{L}_{45}^{1}.$
Remark 6.
The algebras listed below form Lie triple systems.
$\mathcal{L}_{02},\ \mathcal{L}_{06},\ \mathcal{L}_{10},\ \mathcal{L}_{16}^{\alpha},\ \mathcal{L}_{31},\ \mathcal{L}_{35},\ \mathcal{L}_{50}.$
Definition 7 (see [12]).
A Bol algebra is a vector space $\mathfrak{B}$ with a bilinear multiplication [,][-,-][ - , - ] and a trilinear multiplication [,,][-,-,-][ - , - , - ] satisfying:
$\begin{array}[]{ll}({\rm B1})&[x,y]+[y,x]=0,\\ ({\rm B2})&[x,y,z]+[y,x,z]=0,\\ ({\rm B3})&[x,y,z]+[y,z,x]+[z,x,y]=0,\\ ({\rm B4})&[u,v,[x,y,z]]-[[u,v,x],y,z]-[x,[u,v,y],z]-[x,y,[u,v,z]]=0,\\ ({\rm B5})&[[x,y,z],t]-[[x,y,t],z]+[z,t,[x,y]]-[x,y,[z,t]]+[[x,y],[z,t]]=0.\\ \end{array}$
[…]
Theorem 8 (see [2]).
Let $\mathfrak{B}$ be a 4-dimensional complex nilpotent Bol algebra. Then $\mathfrak{B}$ is isomorphic to one of these Lie-Yamaguti algebras $\mathcal{L}_{01},\dots,\mathcal{L}_{41},$ $\mathcal{L}_{45}^{\alpha},\dots,\mathcal{L}_{52}$ or one of the following non-mutually isomorphic algebras:
$\mathfrak{B}_{01}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{2},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{1}]=-\frac{1}{2}e_{4},$
$[e_{1},e_{3},e_{2}]=\frac{1}{2}e_{4},$ $[e_{2},e_{3},e_{2}]=e_{4}.$
$\mathfrak{B}_{02}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{1}]=-\frac{1}{2}e_{4},$
$[e_{1},e_{3},e_{2}]=\frac{1}{2}e_{4},$ $[e_{2},e_{3},e_{2}]=e_{4}.$
$\mathfrak{B}_{03}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{2},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{1}]=-\frac{1}{2}e_{4},$
$[e_{1},e_{3},e_{2}]=\frac{1}{2}e_{4},$ $[e_{2},e_{3},e_{2}]=e_{4}.$
$\mathfrak{B}_{04}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{2},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{1}]=(\alpha-1)e_{4},$ $[e_{1},e_{3},e_{2}]=\alpha e_{4}.$
$\mathfrak{B}_{05}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{3}]=e_{4},$
$[e_{2},e_{3},e_{1}]=(\alpha-1)e_{4},$ $[e_{1},e_{3},e_{2}]=\alpha e_{4}.$
$\mathfrak{B}_{06}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=e_{4},$ $[e_{2},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{3}]=e_{4},$
$[e_{2},e_{3},e_{1}]=(\alpha-1)e_{4},$ $[e_{1},e_{3},e_{2}]=\alpha e_{4}.$
$\mathfrak{B}_{07}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=\alpha e_{4},$ $[e_{1},e_{2},e_{1}]=e_{4},$ $[e_{1},e_{2},e_{3}]=e_{4},$
$[e_{2},e_{3},e_{1}]=e_{4},$ $[e_{1},e_{3},e_{2}]=2e_{4}.$
$\mathfrak{B}_{08}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=\alpha e_{4},$ $[e_{2},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{4},$
$[e_{1},e_{2},e_{3}]=e_{4},$ $[e_{2},e_{3},e_{1}]=e_{4},$ $[e_{1},e_{3},e_{2}]=2e_{4}.$
$\mathfrak{B}_{09}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{2},e_{1}]=e_{3},$ $[e_{1},e_{2},e_{3}]=e_{4},$ $[e_{1},e_{3},e_{2}]=e_{4}.$
$\mathfrak{B}_{10}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{1},e_{3}]=e_{4},$ $[e_{1},e_{2},e_{1}]=e_{3},$
$[e_{1},e_{2},e_{3}]=e_{4},$ $[e_{1},e_{3},e_{2}]=e_{4}.$
Definition 9.
[…]
$\{[x,y],z\}+\{[y,z],x\}+\{[z,x],y\}+[\{x,y\},z]+[\{y,z\},x]+[\{z,x\},y]=0.$
Theorem 10 (see [29]).
Let $\mathfrak{g}$ be a complex $3$-dimensional nilpotent compatible Lie algebra. Then $\mathfrak{g}$ is isomorphic to one of the following pairwise non-isomorphic algebras:
$\mathfrak{L}_{1}$ : $\{e_{1},e_{2}\}=e_{3}.$
$\mathfrak{L}_{2}^{\alpha}$ : $[e_{1},e_{2}]=e_{3},$$\{e_{1},e_{2}\}=\alpha e_{3}.$
Theorem 11 (see [29]).
Let $\mathfrak{g}$ be a complex $4$-dimensional nilpotent compatible Lie algebra. Then $\mathfrak{g}$ is isomorphic to one of the following pairwise non-isomorphic algebras:
${L}_{01}$ : $\{e_{1},e_{2}\}=e_{3}.$
${L}_{02}^{\alpha}$ : $[e_{1},e_{2}]=e_{3},$$\{e_{1},e_{2}\}=\alpha e_{3}.$
${L}_{03}$ :::: $[e_{2},e_{3}]=e_{4},$$\{e_{1},e_{3}\}=e_{4}.$
${L}_{04}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{2},e_{3}]=e_{4},$$\{e_{1},e_{2}\}=e_{4}.$
${L}_{05}^{\alpha}$ : $[e_{2},e_{3}]=\alpha e_{4},$$\{e_{1},e_{2}\}=e_{3},$$\{e_{2},e_{3}\}=e_{4}.$
${L}_{06}$ : $[e_{1},e_{3}]=e_{4},$$\{e_{1},e_{2}\}=e_{3},$$\{e_{2},e_{3}\}=e_{4}.$
${L}_{07}$ :::: $[e_{1},e_{2}]=e_{4},$$\{e_{1},e_{2}\}=e_{3},$$\{e_{2},e_{3}\}=e_{4}.$
${L}_{08}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{3},$$\{e_{1},e_{2}\}=\alpha e_{3},$$\{e_{2},e_{3}\}=e_{4}.$
${L}_{09}^{\alpha,\beta}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{2},e_{3}]=e_{4},$$\{e_{1},e_{2}\}=\alpha e_{3},$$\{e_{2},e_{3}\}=\beta e_{4}.$
${L}_{10}^{\alpha}$ :::: $[e_{1},e_{2}]=e_{3},$ $[e_{2},e_{3}]=e_{4},$$\{e_{1},e_{2}\}=\alpha e_{3},$$\{e_{1},e_{3}\}=e_{4}.$
2 The geometric classification
2.1 Preliminaries: geometric classification
Given a complex vector space ${\mathbb{V}}$ of dimension $n$, the set of bilinear and triliners maps
$\textrm{Bil}({\mathbb{V}}\times{\mathbb{V}},{\mathbb{V}})$cong $\textrm{Hom}({\mathbb{V}}^{\otimes 2},{\mathbb{V}})$cong $({\mathbb{V}}^{*})^{\otimes 2}\otimes{\mathbb{V}},$
$\textrm{Tril}({\mathbb{V}}\times{\mathbb{V}}\times{\mathbb{V}},{\mathbb{V}})$cong $\textrm{Hom}({\mathbb{V}}^{\otimes 3},{\mathbb{V}})$cong $({\mathbb{V}}^{*})^{\otimes 3}\otimes{\mathbb{V}}$
are vector spaces of dimension $n^{3}$ and $n^{4}$, respectively. The set of pairs of bilinear and triliners maps
$\textrm{Bil}({\mathbb{V}}\times{\mathbb{V}},{\mathbb{V}})\oplus\textrm{Tril}({\mathbb{V}}\times{\mathbb{V}}\times{\mathbb{V}},{\mathbb{V}})\cong(({\mathbb{V}}^{*})^{\otimes 2}\otimes{\mathbb{V}})\oplus(({\mathbb{V}}^{*})^{\otimes 3}\otimes{\mathbb{V}})$
which is a vector space of dimension $n^{4}+n^{3}$. This vector space has the structure of the affine space $\mathbb{C}^{n^{4}+n^{3}}$ in the following sense: fixed a basis $e_{1},\ldots,e_{n}$ of ${\mathbb{V}}$, then any pair $(\mu,\mu^{\prime})\in\textrm{Bil}({\mathbb{V}}\times{\mathbb{V}},{\mathbb{V}})\oplus\textrm{Tril}({\mathbb{V}}\times{\mathbb{V}}\times{\mathbb{V}},{\mathbb{V}})$ is determined by some parameters $c_{i,j}^{k},c_{i,j,k}^{l}\in\mathbb{C}$, called structural constants, such that
$\mu(e_{i},e_{j})=\sum_{k=1}^{n}c_{i,j}^{k}e_{k}\textrm{ and }\mu^{\prime}(e_{i},e_{j},e_{k})=\sum_{l=1}^{n}c_{i,j,k}^{l}e_{t}$
which corresponds to a point in the affine space $\mathbb{C}^{n^{4}+n^{3}}$. Then a subset $\mathcal{S}$ of $\textrm{Bil}({\mathbb{V}}\times{\mathbb{V}},{\mathbb{V}})\oplus\textrm{Tril}({\mathbb{V}}\times{\mathbb{V}}\times{\mathbb{V}},{\mathbb{V}})$ corresponds to an algebraic variety, i.e., a Zariski closed set, if there are some polynomial equations in variables $c_{i,j}^{k},c_{i,j,k}^{l}$ with zero locus equal to the set of structural constants of the pairs in $\mathcal{S}$.
Given the identities defining Lie-Yamaguti (Bol) algebras, we can obtain a set of polynomial equations in the variables $c_{i,j}^{k},c_{i,j,k}^{l}$. This class of $n$-dimensional Lie-Yamaguti (Bol) algebras is a variety. Denote it by $\mathcal{T}_{n}$. Now, consider the following action of $\rm{GL}({\mathbb{V}})$ on ${\mathcal{T}}_{n}$:
$(g*\mu)(x,y):=g\mu(g^{-1}x,g^{-1}y),\quad(g*\mu^{\prime})(x,y,z):=g\mu^{\prime}(g^{-1}x,g^{-1}y,g^{-1}z)$
for $g\in\rm{GL}({\mathbb{V}})$, $(\mu,\mu^{\prime})\in\mathcal{T}_{n}$ and for any $x,y,z\in{\mathbb{V}}$. Observe that the $\textrm{GL}({\mathbb{V}})$-orbit of $(\mu,\mu^{\prime})$, denoted $O((\mu,\mu^{\prime}))$, contains all the structural constants of the pairs isomorphic to the Lie-Yamaguti (Bol) algebra with structural constants $(\mu,\mu^{\prime})$.
A geometric classification of a variety of algebras consists of describing the irreducible components of the variety. Recall that any affine variety can be represented uniquely as a finite union of its irreducible components. Note that describing the irreducible components of ${\mathcal{T}_{n}}$ gives us the rigid algebras of the variety, which are those bilinear pairs with an open $\textrm{GL}(\mathbb{V})$-orbit. This is because a bilinear pair is rigid in a variety if and only if the closure of its orbit is an irreducible component of the variety. For this reason, the following notion is convenient. Denote by $\overline{O((\mu,\mu^{\prime}))}$ the closure of the orbit of $(\mu,\mu^{\prime})\in{\mathcal{T}_{n}}$.
Definition 12.
Let ${\rm T}$ and ${\rm T}^{\prime}$ be two $n$-dimensional Lie-Yamaguti (Bol) algebras corresponding to the variety $\mathcal{T}_{n}$ and $(\mu,\mu^{\prime}),(\lambda,\lambda^{\prime})\in\mathcal{T}_{n}$ be their representatives in the affine space, respectively. The algebra ${\rm T}$ is said to degenerate to ${\rm T}^{\prime}$, and we write ${\rm T}\to{\rm T}^{\prime}$, if $(\lambda,\lambda^{\prime})\in\overline{O((\mu,\mu^{\prime}))}$. If ${\rm T}\not\cong{\rm T}^{\prime}$, then we call it a proper degeneration. Conversely, if $(\lambda,\lambda^{\prime})\not\in\overline{O((\mu,\mu^{\prime}))}$ then we say that ${{\rm T}}$ does not degenerate to ${{\rm T}}^{\prime}$ and we write ${{\rm T}}\not\to{{\rm T}}^{\prime}$.
Furthermore, we have the following notion for a parametric family of algebras.
Definition 13.
Let ${{\rm T}}(*)=\{{{\rm T}}(\alpha):{\alpha\in I}\}$ be a family of $n$-dimensional Lie-Yamaguti (Bol) algebras corresponding to ${\mathcal{T}}_{n}$ and let ${{\rm T}}^{\prime}$ be another $n$-dimensional Lie-Yamaguti (Bol) algebra. Suppose that ${{\rm T}}(\alpha)$ is represented by the structure $(\mu(\alpha),\mu^{\prime}(\alpha))\in{\mathcal{T}}_{n}$ for $\alpha\in I$ and ${{\rm T}}^{\prime}$ is represented by the structure $(\lambda,\lambda^{\prime})\in{\mathcal{T}}_{n}$. We say that the family ${{\rm T}}(*)$ degenerates to ${{\rm T}}^{\prime}$, and write ${{\rm T}}(*)\to{{\rm T}}^{\prime}$, if $(\lambda,\lambda^{\prime})\in\overline{\{O((\mu(\alpha),\mu^{\prime}(\alpha)))\}_{\alpha\in I}}$. Conversely, if $(\lambda,\lambda^{\prime})\not\in\overline{\{O((\mu(\alpha),\mu^{\prime}(\alpha)))\}_{\alpha\in I}}$ then we call it a non-degeneration, and we write ${{\rm T}}(*)\not\to{{\rm T}}^{\prime}$.
Observe that ${\rm T}^{\prime}$ corresponds to an irreducible component of $\mathcal{T}_{n}$ (more precisely, $\overline{{\rm T}^{\prime}}$ is an irreducible component) if and only if ${{\rm T}}\not\to{{\rm T}}^{\prime}$ for any $n$-dimensional Lie-Yamaguti (Bol) algebra ${\rm T}$ and ${{\rm T}(*)}\not\to{{\rm T}}^{\prime}$ for any parametric family of $n$-dimensional Lie-Yamaguti (Bol) algebras ${\rm T}(*)$. In this case, we will use the next ideas to prove that a particular algebra corresponds to an irreducible compotent. Firstly, since $\mathrm{dim}\,O((\mu,\mu^{\prime}))=n^{2}-\mathrm{dim}\,\mathfrak{Der}({\rm T})$, then if ${\rm T}\to{\rm T}^{\prime}$ and ${\rm T}\not\cong{\rm T}^{\prime}$, we have that $\mathrm{dim}\,\mathfrak{Der}({\rm T})<\mathrm{dim}\,\mathfrak{Der}({\rm T}^{\prime})$, where $\mathfrak{Der}({\rm T})$ denotes the Lie algebra of derivations of ${\rm T}$. Secondly, to prove degenerations, let ${{\rm T}}$ and ${{\rm T}}^{\prime}$ be two Lie-Yamaguti (Bol) algebras represented by the structures $(\mu,\mu^{\prime})$ and $(\lambda,\lambda^{\prime})$ from ${{\mathcal{T}}}_{n}$, respectively. Let $c_{i,j}^{k},c_{i,j,k}^{l}$ be the structure constants of $(\lambda,\lambda^{\prime})$ in a basis $e_{1},\dots,e_{n}$ of ${\mathbb{V}}$. If there exist $n^{2}$ maps $a_{i}^{j}(t):\mathbb{C}^{*}\to\mathbb{C}$ such that $E_{i}(t)=\sum_{j=1}^{n}a_{i}^{j}(t)e_{j}$ ($1\leq i\leq n$) form a basis of ${\mathbb{V}}$ for any $t\in\mathbb{C}^{*}$ and the structure constants $c_{i,j}^{k}(t),c_{i,j,k}^{l}(t)$ of $(\mu,\mu^{\prime})$ in the basis $E_{1}(t),\dots,E_{n}(t)$ satisfy $\lim\limits_{t\to 0}c_{i,j}^{k}(t)=c_{i,j}^{k}$ and $\lim\limits_{t\to 0}c_{i,j,k}^{l}(t)=c_{i,j,k}^{l}$, then ${{\rm T}}\to{{\rm T}}^{\prime}$. In this case, $E_{1}(t),\dots,E_{n}(t)$ is called a parametrized basis for ${{\rm T}}\to{{\rm T}}^{\prime}$. If $E_{e_{1}}^{t},E_{e_{2}}^{t},E_{e_{3}}^{t},E_{e_{4}}^{t}$ is a parametric basis for ${\bf A}\to{\bf B}$, then we denote a degeneration by ${\bf A}\xrightarrow{(E_{e_{1}}^{t},E_{e_{2}}^{t},E_{e_{3}}^{t},E_{e_{4}}^{t})}{\bf B}$.
Thirdly, to prove non-degenerations we may use a remark that follows from this lemma, see [4] and the references therein.
Lemma 14.
Consider two Lie-Yamaguti (Bol) algebras ${\rm T}$ and ${\rm T}^{\prime}$. Suppose ${\rm T}\to{\rm T}^{\prime}$. Let ${\mathcal{Z}}$ be a Zariski closed set in ${\mathcal{T}}_{n}$ that is stable by the action of the invertible upper (lower) triangular matrices. Then if there is a representation $(\mu,\mu^{\prime})$ of ${\rm T}$ in ${\mathcal{Z}}$, then there is a representation $(\lambda,\lambda^{\prime})$ of ${\rm T}^{\prime}$ in ${\mathcal{Z}}$.
To apply this lemma, we will give the explicit definition of the appropriate stable Zariski closed ${\mathcal{Z}}$ in terms of the variables $c_{i,j}^{k},c_{i,j,k}^{l}$ in each case. For clarity, we assume by convention that $c_{i,j}^{k}=0$ (resp. $c_{i,j,k}^{l}=0$) if $c_{i,j}^{k}$ (resp. $c_{i,j,k}^{l}$) is not explicitly mentioned on the definition of ${\mathcal{Z}}$.
Remark 15.
Moreover, let ${{\rm T}}$ and ${{\rm T}}^{\prime}$ be two Lie-Yamaguti (Bol) algebras represented by the structures $(\mu,\mu^{\prime})$ and $(\lambda,\lambda^{\prime})$ from ${{\rm T}}_{n}$. Suppose ${\rm T}\to{\rm T}^{\prime}$. Then if $\mu,\lambda$ represents algebras ${\rm T}_{0},{\rm T}^{\prime}_{0}$ in the affine space $\mathbb{C}^{n^{3}}$ and $\mu^{\prime},\lambda^{\prime}$ represents algebras ${\rm T}_{1},{\rm T}^{\prime}_{1}$ in the affine space $\mathbb{C}^{n^{4}}$ of algebras with a single multiplication, respectively, we have ${\rm T}_{0}\to{\rm T}^{\prime}_{0}$ and ${\rm T}_{1}\to{\rm T}^{\prime}_{1}$. So for example, $(0,\mu)$ can not degenerate in $(\lambda,0)$ unless $\lambda=0$.
Fourthly, to prove ${{\rm T}}(*)\to{{\rm T}}^{\prime}$, suppose that ${{\rm T}}(\alpha)$ is represented by the structure $(\mu(\alpha),\mu^{\prime}(\alpha))\in{\mathcal{T}}_{n}$ for $\alpha\in I$ and ${{\rm T}}^{\prime}$ is represented by the structure $(\lambda,\lambda^{\prime})\in{\mathcal{T}}_{n}$. Let $c_{i,j}^{k},c_{i,j,k}^{l}$ be the structure constants of $(\lambda,\lambda^{\prime})$ in a basis $e_{1},\dots,e_{n}$ of ${\mathbb{V}}$. If there is a pair of maps $(f,(a_{i}^{j}))$, where $f:\mathbb{C}^{*}\to I$ and $a_{i}^{j}:\mathbb{C}^{*}\to\mathbb{C}$ are such that $E_{i}(t)=\sum_{j=1}^{n}a_{i}^{j}(t)e_{j}$ ($1\leq i\leq n$) form a basis of ${\mathbb{V}}$ for any $t\in\mathbb{C}^{*}$ and the structure constants $c_{i,j}^{k}(t),c_{i,j,k}^{l}(t)$ of $(\mu\big{(}f(t)\big{)},\mu^{\prime}\big{(}f(t)\big{)})$ in the basis $E_{1}(t),\dots,E_{n}(t)$ satisfy $\lim\limits_{t\to 0}c_{i,j}^{k}(t)=c_{i,j}^{k}$ and $\lim\limits_{t\to 0}c_{i,j,k}^{l}(t)=c_{i,j,k}^{l}$, then ${{\rm T}}(*)\to{{\rm T}}^{\prime}$. In this case $E_{1}(t),\dots,E_{n}(t)$ and $f(t)$ are called a parametrized basis and a parametrized index for ${{\rm T}}(*)\to{{\rm T}}^{\prime}$, respectively. Fithly, to prove ${{\rm T}}(*)\not\to{{\rm T}}^{\prime}$, we may use an analogous of Remark 15 for parametric families that follows from Lemma 16, see [4].
Lemma 16.
Consider the family of Lie-Yamaguti (Bol) algebras ${\rm T}(*)$ and the Lie-Yamaguti (Bol) algebra ${\rm T}^{\prime}$. Suppose ${\rm T}(*)\to{\rm T}^{\prime}$. Let ${\mathcal{Z}}$ be a Zariski closed in $\mathcal{T}_{n}$ that is stable by the action of the invertible upper (lower) triangular matrices. Then if there is a representation $(\mu(\alpha),\mu^{\prime}(\alpha))$ of ${\rm T}(\alpha)$ in ${\mathcal{Z}}$ for every $\alpha\in I$, then there is a representation $(\lambda,\lambda^{\prime})$ of ${\rm T}^{\prime}$ in ${\mathcal{Z}}$.
Finally, the following remark simplifies the geometric problem.
Remark 17.
Let $(\mu,\mu^{\prime})$ and $(\lambda,\lambda^{\prime})$ represent two Lie-Yamaguti algebras. Suppose $(\lambda,0)\not\in\overline{O((\mu,0))}$, ((((resp. $(0,\lambda^{\prime})\not\in\overline{O((0,\mu^{\prime}))}$)))), then $(\lambda,\lambda^{\prime})\not\in\overline{O((\mu,\mu^{\prime}))}$. As we construct the classification of Lie-Yamaguti (Bol) algebras from a certain class of algebras with a single multiplication which remains unchanged, this remark becomes very useful.
2.2 The geometric classification of Lie-Yamaguti algebras
In this subsection, we determine all the irreducible components of the variety of four-dimensional nilpotent Lie-Yamaguti algebras.
Theorem 18.
The variety of four-dimensional nilpotent Lie-Yamaguti algebras has dimension $14$ and it has $3$ irreducible components defined by
$\mathcal{C}_{1}=\overline{O({\mathcal{L}}_{19}^{\alpha})},$  $\mathcal{C}_{2}=\overline{O({\mathcal{L}}_{29})},$  and  $\mathcal{C}_{3}=\overline{O({\mathcal{L}}_{44}^{\alpha})}.$
In particular, ${\mathcal{L}}_{29}$ is a rigid algebra.
Proof.
After carefully checking the dimensions of orbit closures of the more important for us algebras, we have
$\dim O({\mathcal{L}}_{19}^{\alpha})=\dim O({\mathcal{L}}_{29})=\dim O({\mathcal{L}}_{44}^{\alpha})=14.$
The following degenerations are observed:
$\mathcal{L}_{44}^{0}$ $\xrightarrow{(te_{1},e_{2},te_{3}+te_{4},e_{4})}$ $\mathcal{L}_{01}$
$\mathcal{L}_{44}^{0}$ $\xrightarrow{(t^{-1}e_{1},t^{3}e_{2},te_{3}+te_{4},e_{4})}$ $\mathcal{L}_{02}$
$\mathcal{L}_{44}^{0}$ $\xrightarrow{(e_{1},te_{2},te_{3}+te_{4},e_{4})}$ $\mathcal{L}_{03}$
$\mathcal{L}_{15}$ $\xrightarrow{(t^{2}e_{1},te_{2},e_{3},t^{2}e_{4})}$ $\mathcal{L}_{04}$
$\mathcal{L}_{15}$ $\xrightarrow{(t^{3}e_{1},te_{2}+e_{3},t(t+1)e_{3},t^{3}(t+1)e_{4})}$ $\mathcal{L}_{05}$
$\mathcal{L}_{07}$ $\xrightarrow{(t^{2}e_{1},te_{2},e_{3},t^{2}e_{4})}$ $\mathcal{L}_{06}$
$\mathcal{L}_{19}^{0}$ $\xrightarrow{(t^{3}e_{1},-t^{2}e_{1}-2t^{2}e_{3},-te_{1}+te_{2}-te_{3},-2t^{5}e_{4})}$ $\mathcal{L}_{07}$
$\mathcal{L}_{19}^{2}$ $\xrightarrow{(-\frac{5t^{2}}{64}e_{1},\frac{t}{16}e_{1}+te_{3},-\frac{1}{4}e_{1}+\frac{1}{3}e_{2}+e_{3},-\frac{5t^{2}}{48}e_{4})}$ $\mathcal{L}_{08}$
$\mathcal{L}_{19}^{-1}$ $\xrightarrow{(\frac{2t^{2}}{1-2t}e_{1},\frac{2t}{1-2t}e_{1}+te_{3},\frac{1}{1-2t}e_{1}-e_{2}+e_{3},\frac{2t^{2}}{2t-1}e_{4})}$ $\mathcal{L}_{09}$
$\mathcal{L}_{15}$ $\xrightarrow{(t^{-1}e_{1},t^{-1}e_{2},t^{-1}e_{3},t^{-3}e_{4})}$ $\mathcal{L}_{10}$
$\mathcal{L}_{15}$ $\xrightarrow{(e_{1},e_{2},t^{-1}e_{3},t^{-1}e_{4})}$ $\mathcal{L}_{11}$
$\mathcal{L}_{15}$ $\xrightarrow{(-\frac{1}{2}(t+2)e_{2},\frac{(t+1)(t+2)^{2}}{t^{2}}e_{1}+\frac{t^{2}+3t+2}{t}e_{2},-\frac{t+2}{2t}e_{3},-\frac{(t+1)(t+2)^{3}}{2t^{3}}e_{4})}$ $\mathcal{L}_{12}$
$\mathcal{L}_{15}$ $\xrightarrow{(t^{-1}e_{1},t^{-1}e_{2},e_{3},t^{-2}e_{4})}$ $\mathcal{L}_{13}$
$\mathcal{L}_{15}$ $\xrightarrow{(-\frac{1}{2}(t+2)e_{2},\frac{(t+2)^{2}}{t^{2}}e_{1}-\frac{t+2}{t}e_{2},\frac{t+2}{2}e_{3},\frac{(t+2)^{3}}{2t^{2}}e_{4})}$ $\mathcal{L}_{14}$
$\mathcal{L}_{19}^{1+t}$ $\xrightarrow{(-\frac{2t}{(2+t)^{2}}e_{1}-\frac{2t}{(2+t)^{2}}e_{2},-\frac{2}{2+t}e_{1}+\frac{2}{2+t}e_{2},\frac{2}{2+t}e_{3},-\frac{8t}{(2+t)^{3}}e_{4})}$ $\mathcal{L}_{15}$
$\mathcal{L}_{19}^{\alpha}$ $\xrightarrow{(t^{-1}e_{1},t^{-1}e_{2},t^{-1}e_{3},t^{-3}e_{4})}$ $\mathcal{L}_{16}^{\alpha}$
$\mathcal{L}_{19}^{\alpha}$ $\xrightarrow{(t^{-1}e_{1},t^{-1}e_{2},e_{3},t^{-2}e_{4})}$ $\mathcal{L}_{17}^{\alpha}$
$\mathcal{L}_{19}^{\alpha}$ $\xrightarrow{(t^{-1}e_{1},e_{2},e_{3},t^{-1}e_{4})}$ $\mathcal{L}_{18}^{\alpha}$
$\mathcal{L}_{44}^{\frac{1}{t^{2}-\alpha}}$ $\xrightarrow{(te_{1},\frac{t^{2}}{t^{2}-\alpha}e_{2},\frac{t^{3}}{t^{2}-\alpha}e_{3}-\frac{\alpha t^{3}}{(t^{2}-\alpha)^{2}}e_{4},\frac{t^{4}}{(t^{2}-\alpha)^{2}}e_{4})}$ $\mathcal{L}_{20}^{\alpha}$
$\mathcal{L}_{44}^{0}$ $\xrightarrow{(te_{1},-te_{2},-t^{2}e_{3}-t^{2}e_{4},t^{3}e_{4})}$ $\mathcal{L}_{21}$
$\mathcal{L}_{23}^{0}$ $\xrightarrow{(te_{1},e_{2},te_{3},te_{4})}$ $\mathcal{L}_{22}$
$\mathcal{L}_{29}$ $\xrightarrow{(-\frac{2t^{2}}{\alpha}e_{1}-\frac{2t^{2}}{\alpha^{2}}e_{2},-\frac{t}{\alpha}e_{1}+\frac{t}{\alpha}e_{2}+\frac{2t}{\alpha^{2}}e_{3},-\frac{2t^{3}}{\alpha^{2}}e_{3}-\frac{4t^{3}}{\alpha^{3}}e_{4},\frac{4t^{5}}{\alpha^{4}}e_{4})}$ $\mathcal{L}_{23}^{\alpha}$
$\mathcal{L}_{23}^{t}$ $\xrightarrow{(e_{1}+t^{-1}e_{3},t^{-1}e_{1}+e_{2},e_{3}-t^{-1}e_{4},e_{4})}$ $\mathcal{L}_{24}$
$\mathcal{L}_{23}^{t^{-1}}$ $\xrightarrow{(te_{1},e_{2},te_{3},te_{4})}$ $\mathcal{L}_{25}$
$\mathcal{L}_{23}^{1}$ $\xrightarrow{(te_{1}+te_{3},e_{1}+e_{2},te_{3}-te_{4},te_{4})}$ $\mathcal{L}_{26}$
$\mathcal{L}_{29}$ $\xrightarrow{(t^{-1}e_{1},t^{-1}e_{2},t^{-2}e_{3},t^{-4}e_{4})}$ $\mathcal{L}_{27}$
$\mathcal{L}_{29}$ $\xrightarrow{(t^{-1}e_{1},e_{2},t^{-1}e_{3},t^{-2}e_{4})}$ $\mathcal{L}_{28}$
$\mathcal{L}_{44}^{1}$ $\xrightarrow{(t^{-1}e_{1},te_{2},t^{-1}e_{3},t^{-2}e_{4})}$ $\mathcal{L}_{30}$
$\mathcal{L}_{44}^{0}$ $\xrightarrow{(t^{-1}e_{1},e_{2},t^{-2}e_{3},-t^{-3}e_{4})}$ $\mathcal{L}_{31}$
$\mathcal{L}_{44}^{0}$ $\xrightarrow{(t^{-1}e_{1},-t^{2}e_{2},-e_{3},-te_{4})}$ $\mathcal{L}_{32}$
$\mathcal{L}_{44}^{-t}$ $\xrightarrow{(t^{-1}e_{1},te_{2},t^{-1}e_{3},-t^{-1}e_{4})}$ $\mathcal{L}_{33}$
$\mathcal{L}_{44}^{t^{2}}$ $\xrightarrow{(t^{-1}e_{1},-t^{2}e_{2},-e_{3},-te_{4})}$ $\mathcal{L}_{34}$
$\mathcal{L}_{44}^{2t^{3}}$ $\xrightarrow{(t^{-1}e_{1}-2t^{2}e_{2},-2t^{3}e_{2}-e_{3},-2te_{3}-2te_{4},-4t^{2}e_{4})}$ $\mathcal{L}_{35}$
$\mathcal{L}_{44}^{\frac{2t^{4}+2t^{3}}{t-1}}$ $\xrightarrow{(\frac{1}{t}e_{1}+\frac{2t^{3}}{1-t}e_{2},\frac{2t^{4}}{1-t}e_{2}+\frac{t}{1-t}e_{3},\frac{2t^{2}}{1-t}e_{3}-\frac{2t^{3}}{(1-t)^{2}}e_{4},-\frac{4t^{4}}{(1-t)^{2}}e_{4})}$ $\mathcal{L}_{36}$
$\mathcal{L}_{44}^{-10t^{3}}$ $\xrightarrow{(t^{-1}e_{1}+2t^{2}e_{2},2t^{3}e_{2}+e_{3},2te_{3}-2te_{4},-4t^{2}e_{4})}$ $\mathcal{L}_{37}$
$\mathcal{L}_{44}^{-t^{2}}$ $\xrightarrow{(t^{-1}e_{1}+te_{2},t^{2}e_{2},e_{3},-e_{4})}$ $\mathcal{L}_{38}$
$\mathcal{L}_{44}^{t^{3}}$ $\xrightarrow{(t^{-1}e_{1}-t^{2}e_{2},-t^{3}e_{2},-te_{3},-t^{2}e_{4})}$ $\mathcal{L}_{39}$
$\mathcal{L}_{44}^{1+t}$ $\xrightarrow{(t^{-1}e_{1}-e_{2},-te_{2},-t^{-1}e_{3},-t^{-2}e_{4})}$ $\mathcal{L}_{40}$
$\mathcal{L}_{44}^{-2t^{3}}$ $\xrightarrow{(t^{-1}e_{1},2t^{4}e_{2}+te_{3},2t^{2}e_{3},-4t^{4}e_{4})}$ $\mathcal{L}_{41}$
$\mathcal{L}_{44}^{0}$ $\xrightarrow{(e_{1},t^{-1}e_{2},t^{-1}e_{3},t^{-2}e_{4})}$ $\mathcal{L}_{42}$
$\mathcal{L}_{44}^{t^{-1}}$ $\xrightarrow{(e_{1},t^{-1}e_{2},t^{-1}e_{3},t^{-2}e_{4})}$ $\mathcal{L}_{43}$
$\mathcal{L}_{44}^{-\frac{1+\alpha}{t}}$ $\xrightarrow{(e_{1}+\frac{\alpha}{t}e_{2},e_{2},e_{3},-t^{-1}e_{4})}$ $\mathcal{L}_{45}^{\alpha}$
$\mathcal{L}_{44}^{\frac{2}{\alpha}}$ $\xrightarrow{(e_{1}-\frac{1}{\alpha}e_{2},-\frac{t}{\alpha}e_{2},-\frac{t}{\alpha}e_{3},-\frac{t}{\alpha^{2}}e_{4})}$ $\mathcal{L}_{46}^{\alpha\neq 0}$
$\mathcal{L}_{44}^{\frac{2(\alpha+1)t}{\alpha-1}}$ $\xrightarrow{(e_{1}-\frac{2\alpha t}{\alpha-1}e_{2},\frac{2t^{2}}{1-\alpha}e_{2}+\frac{t}{1-\alpha}e_{3},\frac{2t^{2}}{1-\alpha}e_{3}-\frac{2\alpha t^{2}}{(1-\alpha)^{2}}e_{4},-\frac{4t^{3}}{(1-\alpha)^{2}}e_{4})}$ $\mathcal{L}_{47}^{\alpha\neq 1}$
$\mathcal{L}_{44}^{-t^{-1}}$ $\xrightarrow{(e_{1}+t^{-1}e_{2},e_{2},e_{3},-t^{-1}e_{4})}$ $\mathcal{L}_{48}$
$\mathcal{L}_{44}^{2t}$ $\xrightarrow{(e_{1}-2te_{2}+e_{3},-2t^{2}e_{2}-te_{3},-2t^{2}e_{3},-4t^{3}e_{4})}$ $\mathcal{L}_{49}$
$\mathcal{L}_{44}^{0}$ $\xrightarrow{(t^{-1}e_{1},2t^{2}e_{2}+t^{-1}e_{3},2e_{3},-4e_{4})}$ $\mathcal{L}_{50}$
$\mathcal{L}_{44}^{0}$ $\xrightarrow{(t^{-1}e_{1},t^{3}e_{2},te_{3},t^{2}e_{4})}$ $\mathcal{L}_{51}$
$\mathcal{L}_{44}^{0}$ $\xrightarrow{(e_{1},2e_{2}+t^{-1}e_{3},2e_{3},-4t^{-1}e_{4})}$ $\mathcal{L}_{52}$
2.3 The geometric classification of Bol algebras
In this section, we determine all the irreducible components of the variety of four-dimensional nilpotent Bol algebras.
Theorem 19.
The variety of four-dimensional nilpotent Bol algebras has dimension $15$ and it has $2$ irreducible components defined by
$\mathcal{C}_{1}=\overline{O({\mathfrak{B}}_{06}^{\alpha})}$  and  $\mathcal{C}_{2}=\overline{O({\mathfrak{B}}_{10})}.$
In particular, ${\mathfrak{B}}_{10}$ is a rigid algebra.
Proof.
From the proof of Theorem 18, we obtain that there are degeneration from the algebras $\mathcal{L}_{19}^{\alpha}$ and $\mathcal{L}_{29}$ to the algebras $\mathcal{L}_{04},$ $\mathcal{L}_{05},$ $\mathcal{L}_{06},$ $\mathcal{L}_{07},$ $\mathcal{L}_{08},$ $\mathcal{L}_{09},$ $\mathcal{L}_{10},$ $\mathcal{L}_{11},$ $\mathcal{L}_{12},$ $\mathcal{L}_{13},$ $\mathcal{L}_{14},$ $\mathcal{L}_{15},$ $\mathcal{L}_{16}^{\alpha},$ $\mathcal{L}_{17}^{\alpha},$ $\mathcal{L}_{18}^{\alpha},$ $\mathcal{L}_{22},$ $\mathcal{L}_{23}^{\alpha},$ $\mathcal{L}_{24},$ $\mathcal{L}_{25},$ $\mathcal{L}_{26},$ $\mathcal{L}_{27},$ $\mathcal{L}_{28}.$
After carefully checking the dimensions of orbit closures of the more important for us algebras, we have
$\dim O({\mathfrak{B}}_{06}^{\alpha})=15$  and  $\dim O({\mathfrak{B}}_{10})=14.$
The following degenerations are observed:
$\mathfrak{B}_{10}$ $\xrightarrow{(te_{1},e_{2},te_{3},e_{4})}$ $\mathcal{L}_{01}$
$\mathfrak{B}_{10}$ $\xrightarrow{(t^{-1}e_{1},t^{4}e_{2},t^{2}e_{3},e_{4})}$ $\mathcal{L}_{02}$
$\mathfrak{B}_{10}$ $\xrightarrow{(e_{1},te_{2},te_{3},e_{4})}$ $\mathcal{L}_{03}$
$\mathfrak{B}_{06}^{\frac{1}{1+\alpha}}$ $\xrightarrow{(-(1+\alpha)e_{1},-(1+\alpha)e_{2},\frac{1}{t}e_{3}+\frac{1}{(1+\alpha)t^{2}}e_{4},-\frac{1+\alpha}{t}e_{4})}$ $\mathcal{L}_{19}^{\alpha\neq-1}$
$\mathfrak{B}_{10}$ $\xrightarrow{(te_{1}+t(\alpha+t^{2})e_{3},te_{2},t^{2}e_{3}+t^{4}e_{4},t^{3}e_{4})}$ $\mathcal{L}_{20}^{\alpha}$
$\mathfrak{B}_{10}$ $\xrightarrow{(te_{2}+te_{3},2te_{1}+2te_{3},-2t^{2}e_{3}-2t^{2}e_{4},-4t^{3}e_{4})}$ $\mathcal{L}_{21}$
$\mathfrak{B}_{06}^{\frac{t+1}{t}}$ $\xrightarrow{(te_{1},te_{2},t^{2}e_{3},t^{3}e_{4})}$ $\mathcal{L}_{29}$
$\mathfrak{B}_{10}$ $\xrightarrow{(t^{-1}e_{1},te_{2},t^{-1}e_{3},t^{-2}e_{4})}$ $\mathcal{L}_{30}$
$\mathfrak{B}_{10}$ $\xrightarrow{(t^{-1}e_{1},t^{-1}e_{2},t^{-3}e_{3},t^{-5}e_{4})}$ $\mathcal{L}_{31}$
$\mathfrak{B}_{10}$ $\xrightarrow{(t^{-1}e_{1}-t^{-4}e_{3},t^{-1}e_{2},t^{-3}e_{3}-t^{-6}e_{4},t^{-5}e_{4})}$ $\mathcal{L}_{32}$
$\mathfrak{B}_{10}$ $\xrightarrow{(t^{-1}e_{1},e_{2},t^{-2}e_{3},t^{-3}e_{4})}$ $\mathcal{L}_{33}$
$\mathfrak{B}_{10}$ $\xrightarrow{(t^{-1}e_{1}-t^{-3}e_{3},e_{2},t^{-2}e_{3}-t^{-4}e_{4},t^{-3}e_{4})}$ $\mathcal{L}_{34}$
$\mathfrak{B}_{10}$ $\xrightarrow{(\frac{1}{t}e_{1}+\frac{1}{t}e_{2},e_{2}+\frac{1}{2t^{3}}e_{3},\frac{1}{t^{2}}e_{3}+\frac{1}{2t^{5}}e_{4},\frac{1}{t^{4}}e_{4})}$ $\mathcal{L}_{35}$
$\mathfrak{B}_{10}$ $\xrightarrow{(\frac{1}{t}e_{1}+e_{2}+\frac{1-t}{2t^{4}}e_{3},te_{2}+\frac{1}{2t^{2}}e_{3},\frac{1}{t}e_{3}+\frac{1}{2t^{4}}e_{4},\frac{1}{t^{2}}e_{4})}$ $\mathcal{L}_{36}$
$\mathfrak{B}_{10}$ $\xrightarrow{(\frac{1}{t}e_{1}+\frac{1}{t}e_{2}-\frac{1}{t^{4}}e_{3},e_{2}+\frac{1}{2t^{3}}e_{3},\frac{1}{t^{2}}e_{3}-\frac{1}{2t^{5}}e_{4},\frac{1}{t^{4}}e_{4})}$ $\mathcal{L}_{37}$
$\mathfrak{B}_{10}$ $\xrightarrow{(\frac{1}{t}e_{1}+\frac{1}{t}e_{2},e_{2},\frac{1}{t^{2}}e_{3},\frac{1}{t^{4}}e_{4})}$ $\mathcal{L}_{38}$
$\mathfrak{B}_{10}$ $\xrightarrow{(\frac{1}{t}e_{1}+\frac{1}{t}e_{2}-\frac{1}{t^{4}}e_{3},e_{2},\frac{1}{t^{2}}e_{3}-\frac{1}{t^{5}}e_{4},\frac{1}{t^{4}}e_{4})}$ $\mathcal{L}_{39}$
$\mathfrak{B}_{10}$ $\xrightarrow{(\frac{1}{t}e_{1}+e_{2},te_{2},\frac{1}{t}e_{3},\frac{1}{t^{2}}e_{4})}$ $\mathcal{L}_{40}$
$\mathfrak{B}_{10}$ $\xrightarrow{(\frac{1}{t}e_{1}+\frac{1}{2t^{4}}e_{3},te_{2}+\frac{1}{2t^{2}}e_{3},\frac{1}{t}e_{3}+\frac{1}{2t^{4}}e_{4},\frac{1}{t^{2}}e_{4})}$ $\mathcal{L}_{41}$
$\mathfrak{B}_{10}$ $\xrightarrow{(e_{1}+\alpha e_{2},te_{2},te_{3},te_{4})}$ $\mathcal{L}_{45}$
$\mathfrak{B}_{10}$ $\xrightarrow{(e_{1}+e_{2}-\alpha e_{3},te_{2},te_{3}-\alpha te_{4},te_{4})}$ $\mathcal{L}_{46}^{\alpha\neq 0}$
$\mathfrak{B}_{10}$ $\xrightarrow{(e_{1}+\alpha e_{2}+\frac{1-\alpha}{2t}e_{3},te_{2}+\frac{1}{2}e_{3},te_{3}+\frac{1}{2}e_{4},te_{4})}$ $\mathcal{L}_{47}$
$\mathfrak{B}_{10}$ $\xrightarrow{(e_{1}+t^{-1}e_{2},te_{2},te_{3},e_{4})}$ $\mathcal{L}_{48}$
$\mathfrak{B}_{10}$ $\xrightarrow{(e_{1}+\frac{1}{t}e_{2}+\frac{t-1}{2t^{2}}e_{3},e_{2}+\frac{1}{2t}e_{3},e_{3}+\frac{1}{2t}e_{4},\frac{1}{t}e_{4})}$ $\mathcal{L}_{49}$
$\mathfrak{B}_{10}$ $\xrightarrow{(\frac{1}{t}e_{1}+\frac{1}{2t^{4}}e_{3},e_{2}+\frac{1}{2t^{3}}e_{3},\frac{1}{t^{2}}e_{3}+\frac{1}{2t^{5}}e_{4},\frac{1}{t^{4}}e_{4})}$ $\mathcal{L}_{50}$
$\mathfrak{B}_{10}$ $\xrightarrow{(e_{1}+t^{-1}e_{3},e_{2}+t^{-1}e_{3},e_{3}+t^{-1}e_{4},2t^{-1}e_{4})}$ $\mathcal{L}_{52}$
$\mathfrak{B}_{10}$ $\xrightarrow{(t^{-1}e_{1}+t^{-4}e_{3},t^{5}e_{2},t^{3}e_{3}+e_{4},-te_{4})}$ $\mathcal{L}_{51}$
$\mathfrak{B}_{02}$ $\xrightarrow{(t^{-1}e_{1},t^{-1}e_{2},t^{-2}e_{3},t^{-4}e_{4})}$ $\mathfrak{B}_{01}$
$\mathfrak{B}_{06}^{\frac{1+2t}{2}}$ $\xrightarrow{(-2te_{2},-e_{1}+e_{2},-2te_{3},4t^{2}e_{4})}$ $\mathfrak{B}_{02}$
$\mathfrak{B}_{02}$ $\xrightarrow{(t^{-1}e_{1},t^{-2}e_{1}+t^{-1}e_{2},t^{-2}e_{3},t^{-4}e_{4})}$ $\mathfrak{B}_{03}$
$\mathfrak{B}_{06}^{\alpha}$ $\xrightarrow{(t^{-1}e_{1},t^{-1}e_{2},t^{-2}e_{3},t^{-4}e_{4})}$ $\mathfrak{B}_{04}^{\alpha}$
$\mathfrak{B}_{06}^{\alpha}$ $\xrightarrow{(t^{-1}e_{1},e_{2},t^{-1}e_{3},t^{-2}e_{4})}$ $\mathfrak{B}_{05}^{\alpha}$
$\mathfrak{B}_{06}^{2+t}$ $\xrightarrow{(\frac{1}{t}e_{1}-\frac{1}{\alpha t^{2}}e_{3},\frac{1}{\alpha}e_{2},\frac{1}{\alpha t}e_{3}+\frac{1}{\alpha^{2}t^{2}}e_{4},\frac{1}{\alpha^{2}t^{2}}e_{4})}$ $\mathfrak{B}_{07}^{\alpha\neq 0}$
$\mathfrak{B}_{06}^{2+t}$ $\xrightarrow{(e_{1}-\frac{1}{\alpha t}e_{3},\frac{1}{\alpha}e_{2},\frac{1}{\alpha}e_{3}+\frac{1}{\alpha^{2}t}e_{4},\frac{1}{\alpha^{2}}e_{4})}$ $\mathfrak{B}_{08}^{\alpha\neq 0}$
$\mathfrak{B}_{10}$ $\xrightarrow{(e_{1},t^{-1}e_{2},t^{-1}e_{3},t^{-2}e_{4})}$ $\mathfrak{B}_{09}$
Below we list all important reasons for necessary non-degeneration.
Non-degenerations reasons
$\mathfrak{B}_{06}^{\alpha}$$\not\rightarrow$ $\mathfrak{B}_{10}$ =$\mathcal{R}=\left\{\begin{array}[]{l}c_{1,2,1}^{3}=0,\ c_{2,3,2}^{4}=0.\end{array}\right\}$
Here $c_{i,j}^{k}$ and $c_{i,j,k}^{l}$ coefficients are structural constants in the basis. ∎
2.4 The geometric classification of compatible Lie algebras
In this subsection, we determine all the irreducible components of the variety of three and four-dimensional nilpotent compatible Lie algebras.
The notions of orbit closure, degeneration, and irreducible components for algebras with two bilinear multiplications are defined analogously to those in Section 2.1; see, for example, [5]. The following results establish that the varieties of three- and four-dimensional nilpotent compatible Lie algebras each consist of a single irreducible component
Proposition 20.
The variety of three-dimensional nilpotent compatible Lie algebras is irreducible and defined by $\mathcal{C}=\overline{O(\mathfrak{L}_{2}^{\alpha})},$ it has dimension $4$.
Proof.
After carefully checking the dimensions of orbit closures of three-dimensional nilpotent compatible Lie algebras, we have $\dim O(\mathfrak{L}_{2}^{\alpha})=4.$ It is not difficult to verify the degeneration $\mathfrak{L}_{2}^{t^{-1}}\ \xrightarrow{(te_{1},e_{2},e_{3})}\ \mathfrak{L}_{1}.$ Therefore, we get that $\mathcal{C}=\overline{O(\mathfrak{L}_{2}^{\alpha})}.$
Theorem 21.
The variety of four-dimensional nilpotent compatible Lie algebras is irreducible and defined by $\mathcal{C}=\overline{O({L_{10}^{\alpha})}},$ it has dimension $13$.
Proof.
After carefully checking the dimensions of orbit closures of the more important for us algebras, we have $\dim O(L_{10}^{\alpha})=13.$
The following degenerations are observed:
$L_{10}^{t^{-1}}$ $\xrightarrow{(e_{1},t^{2}e_{2},te_{3},e_{4})}$ $L_{01}$
$L_{10}^{\alpha}$ $\xrightarrow{(te_{1},e_{2},te_{3},e_{4})}$ $L_{02}^{\alpha}$
$L_{10}^{0}$ $\xrightarrow{(te_{1},te_{2},e_{3},te_{4})}$ $L_{03}$
$L_{10}^{0}$ $\xrightarrow{(te_{1},e_{2},te_{3},te_{4})}$ $L_{04}$
$L_{10}^{t^{-1}}$ $\xrightarrow{(te_{2},e_{1}+\alpha e_{2},-e_{3},-e_{4})}$ $L_{05}^{\alpha}$
$L_{10}^{t^{-1}}$ $\xrightarrow{(e_{2},e_{1},-t^{-1}e_{3},-t^{-1}e_{4})}$ $L_{06}$
$L_{10}^{t^{-1}}$ $\xrightarrow{(t^{2}e_{2},te_{1}-te_{3},-t^{2}e_{3},-t^{3}e_{4})}$ $L_{07}$
$L_{10}^{\alpha}$ $\xrightarrow{(te_{2},e_{1},-te_{3},-te_{4})}$ $L_{08}^{\alpha}$
$L_{10}^{\alpha}$ $\xrightarrow{(te_{2},\beta e_{1}+e_{2},-\beta te_{3},-\beta te_{4})}$ $L_{09}^{\alpha,\beta\neq 0}$
Hence, the variety of four-dimensional nilpotent compatible Lie algebras has a single irreducible component $\overline{O({L_{10}^{\alpha})}}.$